# Learning Massive-scale Partial Correlation Networks in Clinical Multi-omics Studies with HP-ACCORD

### v2

@@ 6.3 Dissecting active regulatory contribution of TFs and repressive epigenomic regulation
Upon understanding the overall structure of the graph, we investigated the regulatory impact of DNA methylation on the expression levels of adjacent genes and dissected the regulatory contribution of TFs on their target gene expression and the repressive effects of hypermethylation in regulatory regions. To this end, we estimated another graph of mRNA expression variables using only the mRNA expression data as input, thereby establishing a reference network which does not exclude the impact of epigenomic regulation. We then compared the resulting mRNA-mRNA network against the one estimated earlier. For the comparison, we removed edges with absolute partial correlation smaller than 0.02 in the network visualization, an arbitrary threshold determined from the histogram of non-zero partial correlation estimates, to minimize the impact of false positive edges as mentioned in §5.
@@
The new mRNA-mRNA network contained 104,599 edges at $\lambda=0.30$ based on epBIC. We observed that the graph estimated without the DNA methylation data, called unadjusted network hereafter, contained 84.5% of the edges from the graph estimated with the DNA methylation data, called adjusted network (Figure 4A). In those edges appearing in both networks, the magnitude of partial correlation estimates in the unadjusted network between mRNA pairs was greater than the one in the adjusted network, as expected (Figure 4B). Overall, the data clearly show that a significant portion of the conditional dependence relationship between two mRNA nodes can be explained away by DNA methylation levels, highlighting the contribution of epigenomic elements in the co-expression patterns.
@@
Since the impact of methylation-driven co-repression of mRNA expression was accounted for, the adjusted network data allows us to infer active co-regulation driven by TFs and co-activators without potential confounding by epignomic regulation. We thus visualized the adjusted network using the Cytoscape software [34] and selected six subnetworks of high connectivity (Figure 4C). We performed hypergeometric probability-based tests for enrichment of biological functions and TFs potentially implicated in each subnetwork of genes. Indeed, the subnetworks showed highly specific enrichment of widely-known cancer associated biological processes, ranging from the extracellular matrix remodeling in the parenchymal environment, immune response and signal transduction, cell cycle and protein translation, and small molecular metabolism with previously well characterized TFs in various cellular processes.
By contrast, the unadjusted network showed a higher degree of connectivity between the nodes (visualization not shown due to lack of legibility) and the overall network lacked clear separation of subnetworks. Interestingly, the network had a large number of nodes with one or two edges connecting to others only. To investigate the quality of the two networks in terms of the recovery of well-validated TF-target relationships, we queried the selected mRNA-mRNA edges in the two networks against bona fide TF-target regulatory network compiled from credible sources [21]. Upon mapping the two mRNA-mRNA networks, 32% and 30% of the edges involved genes encoding a TF or a co-activator in the adjusted and unadjusted networks, respectively. However, while 19% of the reported edges were experimentally validated TF-target pairs in the adjusted network, only 9% of the edges were validated in the unadjusted network. Given the unadjusted network is about 2.6 times larger than the adjusted network, the result implies that the adjusted network recovers the known TF-target regulatory relationships with comparable sensitivity and greater specificity.
@@
In summary, the assessment clearly shows that the network derived from the ultrahigh-dimensional dual-omic data represents more robust relational hypotheses than the network derived from the transcriptomic data alone. With the capability to estimate conditional dependence structure from ultrahigh-dimensional data, spurious associations can be screened out and the results enables subsequent inference of regulatory relationships, which would have been otherwise impossible.
@@ 7 Discussion
In discovery-oriented clinical and basic research applications, technological advances in molecular modalities have steadily increased the number of variables to analyze. Typical dimensionality of data sets exceeds tens or hundreds of thousands, as shown in the motivating dual-omic dataset. The latest technological advances also push sample sizes to thousands, best evidenced by the number of cells analyzed using single-cell resolution omics technologies. When making inference from these truly big data, striking a balance between statistical performance and computational scalability is therefore essential.
The analysis in §6 is a testament to this trade-off. Although the validity of individual edges should be confirmed by experimental means, we verified that the mRNA co-expression network estimated from the ultrahigh-dimensional dual-omic data was largely nested within the network obtained from mRNA expression variables alone, not deviating to a completely distinct result. Moreover, the former network excluding the epigenomic regulation allowed us to recover bona fide TF-target relationships validated in human cells with greater specificity and without loss of sensitivity.
The analysis of dual-omic data presented in this paper would have been impossible without the scalable computation for estimating the underlying graphical model. In an era with new omics modalities added to the multi-omic repertoire and declining cost to produce such data, graphical model estimation in ultrahigh dimensional space will be in increasing demand and computational scalability will thus be as essential as ever. We believe HP-ACCORD paves the way to meeting the demand for graphical model estimation in the modern multi-omic studies and encourages development of related statistical frameworks having scalability in their backbone.

### v3

@@ 6.3 Dissecting active regulatory contribution of TFs and repressive epigenetic regulation
Upon understanding the overall structure of the graph, we next investigated the regulatory impact of DNA methylation on gene expression levels and dissected the regulatory contribution of TFs on their target gene expression from the repressive effects of DNA methylation in the regulatory regions. To this end, we estimated another graph of mRNA expression variables using only the mRNA expression data as input, thereby establishing a reference network which does not exclude the impact of epigenetic regulation. We then compared the resulting mRNA-mRNA network to the one estimated earlier. For the comparison, we removed edges with absolute partial correlation smaller than 0.02 in the network visualization, an arbitrary threshold determined from the histogram of non-zero partial correlation estimates, in order to minimize the impact of false positive edges as mentioned in §5.
@@
The new mRNA-mRNA network contained 104,599 edges at $\lambda=0.30$ based on epBIC. We observed that the graph estimated without the DNA methylation data, called the unadjusted network hereafter, contained 84.5% of the edges from the graph estimated with the DNA methylation data, called the adjusted network (Figure 5A). In the edges appearing in both networks, partial correlation estimates in the unadjusted network were greater in magnitude than those in the adjusted network, as expected (Figure 5B). Overall, the analysis clearly show that a significant portion of the conditional dependence relationship between two mRNA nodes can be explained away by DNA methylation levels, highlighting the contribution of epigenetic elements in the co-expression patterns.
@@
Since the impact of DNA methylation-driven repression of mRNA expression was accounted for, it is reasonable to hypothesize that the adjusted network data allows the analyst to infer active co-regulation driven by TFs and co-activators without the potential confounding by epigenetic regulation. We thus visualized the adjusted network using Cytoscape software [36] and selected six sub-networks of high connectivity (Figure 5C). We performed hypergeometric probability-based test for the enrichment of biological functions and TFs in each subnetwork. The subnetworks showed specific enrichment of cancer-associated biological processes, ranging from extracellular matrix remodeling in the parenchymal environment, immune response and signal transduction, cell cycle and protein translation, and small molecular metabolism with previously well-characterized TFs in various cellular processes. By contrast, the unadjusted network showed a higher degree of connectivity between the nodes (visualization not shown due to lack of legibility), and the overall network lacked clear separation of subnetworks.
Figure 6: Recovery of potentially co-regulated gene pairs sharing one or more common TFs in human cells at a range of $\lambda$ values and at the corresponding network sizes. The curves show the percentages of TF-sharing gene pairs in the selected edges in the DNA methylation adjusted network at $\lambda=0.45,0.48,\ldots,0.66$ (blue) and in the unadjusted network at $\lambda=0.30,0.33,\ldots,0.51$ (red). The large circles correspond to the $\lambda$ values optimized using epBIC in the two networks.
To investigate the quality of the two networks in terms of the recovery of co-regulated genes, we compared how often the reported mRNA-mRNA edges have previously been validated using a compendium of bona fide TF-target regulatory relationships compiled from credible sources. Specifically, we calculated the proportion of edges connecting genes with at least one shared TFs in the validated TF-target relationships [45, 46, 35, 12]. The comparison showed that the adjusted network has greater enrichment of gene pairs sharing common TFs and co-activators (13.6%) than the unadjusted network (9.4%) at the respective optimal $\lambda$ values selected using epBIC (Figure 6, left panel). Considering that the latter network is three times as large as the former, as indicated by large circles in the diagrams (right panel), the results reaffirm that the network derived from the analysis accounting for the impact of DNA methylation confers greater specificity in the reported gene-to-gene connections. When we compared the networks with more stringent selection of edges at equivalent network sizes, the pattern also remained consistent (right panel): the adjusted network captures gene pairs sharing common TFs better than the unadjusted network.
@@
In summary, the evaluation shows that the network derived from the dual-omic data generates more robust relational hypotheses than the network derived from the transcriptomic data alone. By directly estimating conditional dependence structure from ultrahigh-dimensional data, spurious associations can be screened out and the filtered data improves the quality of biological inference of regulatory relationships. It goes without saying that the superior performance in the present case study comes from the scalable computation enabled by the HP-ACCORD framework.
@@ 7 Discussion
In discovery-oriented clinical and molecular biology research, technological advances have steadily increased the number of variables that can be analyzed. Typical data sets have more than tens or hundreds of thousands of variables. The latest technological advances also push sample sizes beyond thousands, best evidenced by the surge in single-cell resolution profiles in the literature. When making biological inferences from these truly big data, striking a balance between statistical performance and computational scalability is therefore essential.
The analysis in §6 is a testament to this trade-off. We acknowledge that the validity of individual edges reported in the analysis should be subjected to experimental confirmation. However, we verified that the mRNA co-expression network estimated from the ultra high-dimensional dual-omic data was largely nested within the network obtained from mRNA expression variables alone, not deviating to a completely distinct result. Moreover, the analysis incorporating epigenomic profiles allowed us to recover potentially co-regulated gene pairs in higher percentage with a smaller network than the latter, through the evaluation for the recovery of bona fide TF-target relationships validated in human cells.
The analysis presented in this paper would have been impossible without the scalable computation. In an era where new omics modalities are constantly being added to the multi-omics repertoire, graphical model estimation in ultra-high-dimensional spaces will become increasingly necessary for data integration, and computational scalability will remain as essential as ever. We believe HP-ACCORD paves the way to meeting this unmet demand and encourages consideration of computational scalability in the development of other statistical frameworks in their backbone.